What is the parabola graph of y^2=10x
Answer by anantha(86)

sol
y^2=10x
dividing by 10 on both sides
y^2/10=10x/10
x=y^2/10
put y=1 then x=y^2/10=1^2/10=1/10=0.1
put y=2 then x=y^2/10=2^2/10=2*2/10=4/10=0.4
put y=3 then x=y^2/10=3^2/10=9/10=0.9
put y=0 then x=y^2/10=0/10=0
put y=-1 then x=y^2/10=(-1)^2/10=1/10=0.1
put y=-2 then x=y^2/10=(-2)^2/10=4/10=0.4
put y=-3 then x=y^2/10=(-3)^2/10=9/10=0.9
the coordinates (x,y) are (0.1,1),(0.4,2),(0.9,3),(0,0),(0.1,-1),(0.4,-2),(0.9,-3)
now draw the graph of the above coordinates using graph paper